# An Information Theory of Finite Abstractions and their Fundamental Scalability Limits

### v1

@@ 6.2 A 3D nonlinear system and abstractions with uniform grids
and $\mathcal{X}=[-1,1]^{3},$ which is forward invariant under $f$. This system has multiple equilibria, hence the origin is not stable in $\mathcal{X}$. For each NN in $\{10,20,50,100\},$ we build abstractions $A_{N}$ by using uniform partitioning of $\mathcal{X}$ with grids of size $N\times N\times N$ and determining the transition map using interval arithmetic. Then, we compute the distortion lower bound from Theorem 5.2 using Prop. 5.3 and Prop. 5.4, case 3.888The entropies were computed using Monte-Carlo integration with 10000 samples, while Jacobians and the Lipschitz constant were determined using automatic differentiation. Furthermore, lower bounds were also computed by picking $c_{{\mathcal{B}}_{l}^{S}}=v_{3}$, in light of Remark 5. The resulting distortion lower bound curves can be seen in Fig. 6. In this case, as the abstraction we construct is not necessarily the optimal one, its expected distortion is generally 100x higher than the fundamental lower bound. Still, this demonstrates the validity of the lower bound, even in cases with nonlinear dynamics; even more importantly, it indicates how conservative standard abstractions with uniform grids might be.
@@ 7 Conclusion and Future Research:
We have developed a statistical, quantitative theory on the accuracy-size tradeoff of finite abstractions of dynamical systems. Through this theory, we have uncovered fundamental limits on their scalability: given the system dynamics, we have obtained a fundamental bound on the achievable abstraction accuracy, for a given abstraction size. To that end, we have established connections with rate-distortion theory. From an information-theoretic perspective, we have developed rate-distortion theory for the particular class of encoder-decoder pairs that abstractions constitute: set-based, with set-based distortion. Overall, this novel theory quantifies scalability limits of abstractions, and provides insights on how the complexity of the dynamics to be abstracted dictates these limits.
Most importantly, the developed theory may be employed to construct minimal abstractions, harnessing their full scalability potential. From this work, it becomes clear that, to construct minimal abstractions, one has to solve the problem of encoding trajectories of dynamical systems, through coverings in a high-dimensional, ambient space. In fact, this has already been demonstrated, in Section 6.1, where we construct a minimal abstraction of the doubling-map dynamics. Future research will thus focus on the general problem of constructing minimal abstractions. Towards that goal, information-theoretic algorithms optimizing the rate-distortion tradeoff, such as the information bottleneck method (see [32]), could be adapted for abstractions.
@@ 8 Technical Results and Proofs
Let $M\subset\mathbb{R}^{n}$ be a finite union of bounded, disjoint, $m$-dimensional $C^{1}$-manifolds. Let $X$ be a random variable in $M$ with probability measure $\mu_{X}\ll\mathcal{H}^{m}_{M}$ and density $p=\frac{\mathrm{d}\mu_{x}}{\mathrm{d}\mathcal{H}^{m}_{M}}$. Then, for any collection $\mathcal{Y}\coloneqq\{Y_{i}\}_{i=1}^{N}$ of NN measurable, $n$-dimensional sets $Y_{i}\subseteq\mathbb{R}^{n}$ covering $M$, the following holds for any $s\in(1,\infty]$:
@@
We make use of Prop. 5.1. Take (8) and minimize both sides over all possible partitions $\mathcal{Y}$ with size $|\mathcal{Y}|\leq e^{R}$ and associated abstractions AA. We have
@@
where recall that $\mathbb{E}_{\xi_{0}}[\cdot]=\mathbb{E}_{\xi}[\cdot]$, and that for a given abstraction AA with corresponding encoder-decoder pair $s_{A},g_{A}$, we have $\xi_{q_{A}}=g_{q_{A}}(s_{q_{A}}(\xi))$ with $s_{q_{A}}(\xi)=g_{A}(s_{A}(\xi))$ and $g_{q_{A}}(z)=x_{c}(z)$, where $x_{c}(z)$ is the Chebyshev center of the set $z$; and $r_{c}(\Omega_{A})$ is the Chebyshev radius of $\Omega_{A}$. Thus, $\xi_{q_{A}}$ is the output of the encoder-decoder pair $(s_{q_{A}},g_{q_{A}})$ with rate $R$ and message $\xi$. Hence, the first term in the left-hand side of the above inequality, can be lower bounded by employing Thm. 2.1, to obtain:
@@
To bound the second term, we employ Lemma 8.1. Notice that the abstraction’s outputs $\Omega_{A}$ are nlnl-dimensional and define a cover [^9] This cover is precisely $\mathcal{Z}:=\{Z:Z=g_{A}(s_{A}(x_{0})),x_{0}\in\mathcal{X}\}$ and note that $s_{A}(x)$ takes values in the set $|\mathcal{Y}|$. Thus $|\mathcal{Z}|=|\mathcal{Y}|$. of ${\mathcal{B}}_{l}^{S}$ (which is $n$-dimensional) with cardinality $|\mathcal{Y}|\leq e^{R}$ (the same as the state-space partition). Thus, the term $\inf_{A,|\mathcal{Y}|\leq e^{R}}\mathbb{E}[r_{c}^{2}(\Omega_{A})]$ can be lower bounded as in (15), where we replace $m$ by $n$, $M$ by ${\mathcal{B}}_{l}^{S}$, NN by $e^{R}$, and $\mu_{X}$ by $\mu_{\xi}$. ∎
Proof of Prop. 5.3.
@@
and, applying (17) gives (12). Finally, in the particular case of $s=\infty,$ we have
@@
which gives the desired result by exploiting the fact that $\log$ is monotonically increasing.
@@
Case (1) follows trivially by the observation that ${\mathcal{B}}_{l}^{S}$ is an $n$-dimensional affine subset of $\mathbb{R}^{nl}$, and that the intersection of an nlnl-ball of radius $r$ and a plane of dimension $n$ is a ball of dimension $n$ and radius $\leq r.$ Hence, $\mathcal{H}_{{\mathcal{B}}_{l}^{S}}(B(z,\delta))\leq v_{n}\delta^{n}$, for all $z\in\mathbb{R}^{nl}$.

### v3

@@ 6.2 A 3D nonlinear system and abstractions with uniform grids
and $\mathcal{X}=[-1,1]^{3},$ which is forward invariant under $f$. This system has multiple equilibria, hence the origin is not stable in $\mathcal{X}$. For each $N$ in $\{10,20,50,100\},$ we build abstractions $A_{N}$ by using uniform partitioning of $\mathcal{X}$ with grids of size $N\times N\times N$ and determining the transition map using interval arithmetic. Then, we compute the distortion lower bound from Theorem 5.2 using Prop. 5.4 and Prop. 5.5, case 3.999The entropies were computed using Monte-Carlo integration with 10000 samples, while Jacobians and the Lipschitz constant were determined using automatic differentiation. Furthermore, lower bounds were also computed by picking $c_{{\mathcal{B}}_{l}^{S}}=v_{3}$, in light of Remark 6. The resulting distortion lower bound curves can be seen in Fig. 6. In this case, as the abstraction we construct is not necessarily the optimal one, its expected distortion is generally 100x higher than the fundamental lower bound. Still, this demonstrates the validity of the lower bound, even in cases with complex, nonlinear dynamics; even more importantly, it indicates how conservative standard abstractions with uniform grids might be.
@@ 7 Conclusion and Future Research:
We have developed a statistical, quantitative theory on the accuracy-size tradeoff of finite abstractions of dynamical systems. Through this theory, we have uncovered fundamental limits on their scalability: given the system dynamics, we have obtained a fundamental bound on the achievable abstraction accuracy, for a given abstraction size; and, conversely, a fundamental lower bound on the abstraction size, for a prescribed accuracy. To that end, we have established connections with rate-distortion theory. From an information-theoretic perspective, we have developed rate-distortion theory for the particular class of encoder-decoder pairs that abstractions constitute: set-based, with set-based distortion. Overall, this novel theory quantifies scalability limits of abstractions, and provides insights on how the complexity of the dynamics to be abstracted dictates these limits.
Most importantly, the developed theory may be employed to construct minimal abstractions, harnessing their full scalability potential. From this work, it becomes clear that, to construct minimal abstractions, one has to solve the problem of encoding trajectories of dynamical systems, through coverings in a high-dimensional ambient space. In fact, this has already been demonstrated, in Section 6.1, where we construct a minimal abstraction of the doubling-map dynamics. Future research will thus focus on the general problem of constructing minimal abstractions. Towards that goal, information-theoretic algorithms optimizing the rate-distortion tradeoff, such as the information bottleneck method [33], could be adapted for abstractions.
@@ 8 Technical Results and Proofs
Let $M\subset\mathbb{R}^{n}$ be a finite union of bounded, disjoint, $m$-dimensional $C^{1}$-manifolds. Let $X$ be a random variable in $M$ with probability measure $\mu_{X}\ll\mathcal{H}^{m}_{M}$ and density $p=\frac{\mathrm{d}\mu_{x}}{\mathrm{d}\mathcal{H}^{m}_{M}}$. Then, for any collection $\mathcal{Y}\coloneqq\{Y_{i}\}_{i=1}^{N}$ of $N$ measurable, $n$-dimensional sets $Y_{i}\subseteq\mathbb{R}^{n}$ covering $M$, the following holds for any $s\in(1,\infty]$:
@@
We make use of Prop. 5.1. Take (7) and minimize both sides over all possible partitions $\mathcal{Y}$ with size $|\mathcal{Y}|\leq e^{R}$ and associated abstractions $A$. We have
@@
where recall that $\mathbb{E}_{\xi_{0}}[\cdot]=\mathbb{E}_{\xi}[\cdot]$, and that for a given abstraction $A$ with corresponding encoder-decoder pair $s_{A},g_{A}$, we have $\xi_{q_{A}}=g_{q_{A}}(s_{q_{A}}(\xi))$ with $s_{q_{A}}(\xi)=g_{A}(s_{A}(\xi))$ and $g_{q_{A}}(z)=x_{c}(z)$, where $x_{c}(z)$ is the Chebyshev center of the set $z$; and $r_{c}(\Omega_{A})$ is the Chebyshev radius of $\Omega_{A}$. Thus, $\xi_{q_{A}}$ is the output of the encoder-decoder pair $(s_{q_{A}},g_{q_{A}})$ with rate $R$ and message $\xi$. Hence, the first term in the left-hand side of the above inequality, can be lower bounded by employing Thm. 2.1, to obtain:
@@
To bound the second term, we employ Lemma 8.1. Notice that the abstraction’s outputs $\Omega_{A}$ are $nl$-dimensional and define a cover [^10] This cover is precisely $\mathcal{Z}:=\{Z:Z=g_{A}(s_{A}(x_{0})),x_{0}\in\mathcal{X}\}$ and note that $s_{A}(x)$ takes values in the set $|\mathcal{Y}|$. Thus $|\mathcal{Z}|=|\mathcal{Y}|$. of ${\mathcal{B}}_{l}^{S}$ (which is $n$-dimensional) with cardinality $|\mathcal{Y}|\leq e^{R}$ (the same as the state-space partition). Thus, the term $\inf_{A,|\mathcal{Y}|\leq e^{R}}\mathbb{E}[r_{c}^{2}(\Omega_{A})]$ can be lower bounded as in (17), where we replace $m$ by $n$, $M$ by ${\mathcal{B}}_{l}^{S}$, $N$ by $e^{R}$, and $\mu_{X}$ by $\mu_{\xi}$. ∎
Proof of Prop. 5.4.
@@
and, applying (19) gives (12). In the particular case of $s=\infty,$ we have
@@
which gives (13) by exploiting the fact that $\log$ is monotonically increasing.
Finally, since $J_{b_{l}}^{\mkern-1.5mu\mathrm{T}}\!J_{b_{l}}=\mathrm{I}+J_{f}^{\mkern-1.5mu\mathrm{T}}\!J_{f}+\cdots\succeq\mathrm{I},$ we have that $\det(J_{b_{l}}^{\mkern-1.5mu\mathrm{T}}\!J_{b_{l}})\geq 1$, establishing that $h(\xi)\geq h(\xi_{0})$ and $h_{s}(\xi)\geq h_{s}(\xi_{0})$ for any $s>1.$
@@
Case (1) follows trivially by the observation that ${\mathcal{B}}_{l}^{S}$ is an $n$-dimensional affine subset of $\mathbb{R}^{nl}$, and that the intersection of an $nl$-ball of radius $r$ and a plane of dimension $n$ is a ball of dimension $n$ and radius $\leq r.$ Hence, $\mathcal{H}_{{\mathcal{B}}_{l}^{S}}(B(z,\delta))\leq v_{n}\delta^{n}$, for all $z\in\mathbb{R}^{nl}$.